1. General status updates/Schedule/Purchases review.
2. Status of DTC to CRV ROC interface.
3. Status of hardware acceptance testing.

Lei - CRV received 25 pilot FEBs to KSU. Two minor issues.

Eric - cut a new version of mu2e_artdaq_core and pcie for Offline builds to help GIani. Kerberos problems at the test stand. Backup today. Requesting a new account for a student by Vivian -- but requested a new host key (by accident).

John -- hoping to have 60 DTC+Timing Card pairs .. hopefully the remainder by next week. Patch panel cassettes had errors with Rx Tx routing .. we sent 36 cassettes back to vendor, Fiber Instruments. Causing vendor documentation and re-testing. Putting 10 complete modules in one Firefly box -- will start going on PREP shelf.

Vivian -- Would like a DTC on daq08, has new student Yunyi (Michigan Tracker with Monica, firmware), hoping he sets up machine in lab 3. Still struggling with new artdaq release. Should checkin DTC/CFO projects into Redmine. Tuesdays at 11:30 Slow Controls/EPICS meeting.

Rick -- Check the link ID to verify patch panels. Possibly use loopback marker. Now DTCs do not need an external CFO once Rick implements the fix (take clock from local crystal rather than jitter attenuator). DTC acceptance test stand is going with 5 machines (daq 12,14,15,16,17), 2 DTCs at a time (2?-3? different mother boards in servers). Eric -- we need specify the mother board model number we want for production purchase. First batch of 10 had no problems. 1 in second batch had regulators down when in PCIe slot. Plan for third batch on Monday.

Bertrand -- Doesn’t thing 100 us is special.

Ron --

Greg -- How do we sync ExtMon/STM/Event in Offline analysis? Investigated how CMS handled this. Found a document (how to use CFO).

Antonio -- First priority is timeout of request to EPICS server, and then improve the widget infrastructure - and then try other plotting libraries.

Glenn --

Ryan -- start thinking about server purchase, Compiling accelerators, Data loggers.